OpenStudy (anonymous):

Differentiate function; G(x)=sin square root of x^2+1

OpenStudy (anonymous):

G'(x) = cos (sqrt(x^2+1) * 1/2(sqrt(x^2+1) * 2x

OpenStudy (anonymous):

\[G'(x) = \cos(\sqrt{x^2+1})*\frac{ x }{ \sqrt{x^2+1} }\]

OpenStudy (anonymous):

@Yahoo! are these the steps

OpenStudy (anonymous):

Use Chain Rule